OpenStudy (anonymous):

LEGEN...waitforit...DARIDDLE: How many people should be in a room so that there is a 99% probability that two or more people share the same birthday? Hint: whoever gets it right with solutions gets a medal

hartnn (hartnn):

the dreaded riddles are back!

OpenStudy (anonymous):

these ones are legendary. i don't make puny riddles like the original

OpenStudy (anonymous):

so @hartnn do you know the answer?

hartnn (hartnn):

haven't tried yet...i would always prefer helping than satisfying my own thirst of solving puzzles...i'll try that in free time :)

OpenStudy (anonymous):

helping huh. haven't tried that in a while

hartnn (hartnn):

i try these problems by taking small numbers first let there be just 2 people in that room then the probability "that two or more people share the same birthday" will be 1- probability that BOTH will not have same birthday (assuming 366 days a year) so, \(\large 1- \dfrac{366}{366}\times \dfrac{365}{366}\) now extending this for 'n' people \(\large 0.99 =1- \dfrac{\dfrac{366!}{(366-n)!}}{366^n}\) am i on right path ? (wondering how the hell can i solve that for n :O O.o)

OpenStudy (anonymous):

lol. am not saying if you're on the right path or not. just tell me the answer and i'll tell you if it's right. that's how riddles work

hartnn (hartnn):

rounding to nearest integer i am getting 55 people (infact 55 or more!)

OpenStudy (anonymous):

ooooh so close

OpenStudy (anonymous):

but no

hartnn (hartnn):

i had my chance, i'll let others try....